# RELATE: Relation Extraction in Biomedical Abstracts with LLMs and Ontology Constraints

### v1

@@ Appendix B SAPBERT Workflow Enhancement
Figure 4: SapBERT Augmented System workflow diagram showing the three-stage pipeline as in Figure LABEL:fig:workflow: (1) Ontology Preprocessing, (2) Similarity-Based Retrieval, and (3) Contextual Refinement. Just as in Figure LABEL:fig:workflow, stage 3.1 changes only if the ontology schema or the LLM and embedding Model changes. However, the one-stage preprocessing of raw ontology predicate data involves two embedding generations- LLM-generated embedding and SapBERT finetuned stored embedding. Also, given a raw input triple with abstract context, the frequently executed stage—similarity-based retrieval and contextual refinement—encompasses the input relation’s dual embedding transformation, dual similarity search on the pair of stored embeddings, merger of top-k candidates from each search, and final contextual reranking.
Appendix C Predicate and Descriptor Distribution
@@
tab:top_predicates_distribution_chemprot Predicate Descriptors Percentage (%) substrate 6 20.00 agonist 5 16.67 upregulator 4 13.33 downregulator 4 13.33 regulator 3 10.00 part of 2 6.67 antagonist 2 6.67 modulator 2 6.67 cofactor 2 6.67
Table 2: Chemprot Dataset Top Predicates by Descriptor Count and Percentage Contribution
@@
tab:top_predicates_distribution_heal Predicate Descriptors Percentage (%) coexists with 178 39.04 related to 96 21.05 located in 40 8.77 part of 34 7.46 temporally related to 30 6.58 has part 21 4.61 precedes 15 3.29 causes 15 3.29 affects 14 3.07 has output 13 2.85
Table 3: Biolink Dataset Top Predicates by Descriptor Count and Percentage Contribution
Table LABEL:tab:top_predicates_distribution_chemprot and Table LABEL:tab:top_predicates_distribution_heal showed the top 10 predicates that account for the majority of descriptors in the dataset. substrate and agonist together accounting for over one-third of all descriptors in Table LABEL:tab:top_predicates_distribution_chemprot while Table LABEL:tab:top_predicates_distribution_heal dominated by predicates such as coexists with and related to.

### v2

@@ Appendix B RELATE Stages Algorithms
Figure 4: SapBERT-augmented RELATE pipeline as in Figure LABEL:fig:workflow. The ontology preprocessing now involves both LLM-generated embeddings and SapBERT finetuned embeddings. Additional changes include dual embedding of the relation text, dual similarity searches on the stored embeddings, and merging of the two top-$k$ candidates.
Appendix C Additional Analysis
C.1 ChemProt Pre/Post Refinement Accuracies: Error Attribution Analysis
Our results revealed an unexpected pattern: Stage 2 (Section 3.2) retrieval achieved 52.8% accuracy@1, but after Stage 3 (Section 3.3) refinement, final accuracy dropped to 46.6% (Table LABEL:tab:chemprot_overall). To understand this decline, we analyzed the 6.2% set failures and attributed them to specific causes.
C.1.1 Stage 3 Impact Analysis
We identified all cases where Stage 3 (Section 3.3) changed Stage 2’s (Section 3.3) @1 prediction and categorized them by outcome:
  1. 1.
    Stage 3 harmful changes (146 cases): Stage 3 changed a correct @1 prediction to an incorrect answer
  2. 2.
    Stage 3 beneficial changes (103 cases): Stage 3 rescued failures by selecting the correct answer from ranks 2-10
The net impact was -43 cases (146 harmful - 103 beneficial), accounting for the observed 6.2% accuracy decline. This analysis focuses on the 146 harmful changes to understand why Stage 3 overrode correct predictions.
C.1.2 Systematic Over-Specification Pattern
Detailed examination of the 146 harmful changes revealed highly concentrated error patterns rather than random failures (Table 2). A single transformation pattern in Stage 2 correctly predicted “downregulator” but Stage 3 changed it to “inhibits”; accounting for 50.7% of all harmful changes. The top five patterns captured 83.6% of errors, indicating systematic bias.
Table 2: Stage 3 Over-Specification Error Patterns with Original Free-Text Relationship Variants
Ground Truth Stage 3 Original Free-Text Variants Count %
downregulator inhibits “inhibits” (68), “inhibits the activity of” (4), “inhibits secretion of” (1), “inhibits the activities of” (1) 74 50.7
downregulator antagonist “blocks” (8), “inhibits” (6), “blockers of” (1), “is a poor inhibitor of” (1), “tyrosine kinase inhibitor” (1), “interferes with inhibitory action on” (1) 18 12.3
downregulator inhibitor “is an inhibitor of” (3), “inhibits” (3), “inhibitor of” (2), “selectively inhibited” (1), “preferentially inhibits” (1), “inhibitory activity” (1), +4 more 15 10.3
upregulator agonist “activates” (4), “induces expression of” (3), “stimulates” (2), “stimulates the transcription of” (1), “increases phosphorylation of” (1) 11 7.5
upregulator activates “activates” (4) 4 2.7
Top 5 patterns 122 83.6%
upregulator modulator “activates” (1), “stimulates activity of” (1), “stimulates the activity of” (1) 3 2.1
downregulator regulator “inhibits” (1), “is abolished by” (1), “reduces concentrations of” (1) 3 2.1
downregulator blocks “blocks” (3) 3 2.1
substrate regulator “impacts metabolism of” (2) 2 1.4
Other patterns (6 additional) 12 8.2
Total 146 100.0%
Analysis of the original free-text relationships revealed the fundamental mechanism:
  1. 1.
    Stage 3 preserved mechanistic specificity from source text rather than maintaining ontology-level abstraction. In the dominant error pattern (downregulator → inhibits), 92% of cases (68/74) had “inhibits” as the original free-text relationship
  2. 2.
    Stage 2 (retrieval-based) correctly abstracted to the ontology-level category “downregulator,” matching ChemProt’s annotation guidelines. However, Stage 3 (LLM-based refinement) reintroduced the mechanistic specificity from the source text, selecting “inhibits” because that was the literal wording in the abstract.
This pattern extended across error types: “blocks” appeared in 61% (11/18) of cases where Stage 3 incorrectly changed “downregulator” to “antagonist,” and “activates” appeared in 73% (8/11) of cases where Stage 3 changed “upregulator” to “agonist.”
C.1.3 Interpretation
These findings illuminate a fundamental tension between retrieval-based and generative approaches to relation extraction. The embedding-based retrieval (Stage 2 (Section 3.2)) successfully learned ontology-level abstractions from the training data, correctly mapping mechanistic phrases to categorical predicates. However, the LLM refinement (Stage 3 (Section 3.2)), despite being trained extensively on biomedical literature, demonstrated a preference for preserving the literal mechanistic terminology found in source text.
The concentration of errors (84%) captured by just five patterns suggests that targeted interventions could recover most of the 6.2% accuracy loss:
  1. 1.
    Confidence-based bypass: Skip Stage 3 when Stage 2’s top-1 confidence exceeds a threshold (e.g., 0.8)
  2. 2.
    Ontology-constrained prompting: Explicitly instruct the LLM to maintain categorical abstraction levels
  3. 3.
    Restricted candidate sets: Only allow Stage 3 to choose from predicates at the same abstraction level as the top-k Stage 2 candidates
C.2 Rejected Relations
RELATE includes a rejection mechanism (NONE option, Section 3.3) to filter relations that cannot be reliably mapped to ontology predicates. Here, we analyze rejection patterns across both datasets.
C.2.1 HEAL Dataset Rejections
In the HEAL dataset of 2,400 abstracts, RELATE rejected 10 relationships (rejection rate: 0.4%). Manual review confirmed that all 10 rejections were appropriate (100% precision), filtering non-biomedical associations.
The rejected relationships fell into distinct categories:
  1. 1.
    Genetic nomenclature specifications (3 cases): Relations describing genetic naming conventions rather than biological relationships (e.g., gene symbol designations)
  2. 2.
    Clinical procedures and documentation (4 cases): Administrative or procedural relationships such as “Clinicians prescribe Opioid” and “Women on BUP had documented Opioid misuse.”
  3. 3.
    Methodological relationships (2 cases): Experimental procedures such as “Anisomycin used to detect GNPTG.”
  4. 4.
    Specialized reporting contexts (1 case): Context-specific descriptions not representing biological relationships.
These rejections demonstrate RELATE’s ability to distinguish between legitimate biomedical relationships and other types of associations commonly found in scientific literature. All rejections were semantically distant from Biolink predicates, making the NONE decision appropriate.
C.2.2 ChemProt Dataset Rejections
Analysis of the ChemProt test set (695 cases with ground truth labels) revealed 18 instances where RELATE selected NONE. Of these, only 5 were correct rejections, yielding a rejection precision of 27.8% (72% false rejection rate among rejected cases).
The 13 incorrect rejections occurred despite appropriate candidate predicates being retrieved with moderate-to-high similarity scores during Stage 2 (Section 3.2). Examples include:
  1. 1.
    “loperamide modifies corticotropin-releasing hormone action” retrieved modulator with similarity score 0.661, yet was rejected in Stage 3.
  2. 2.
    “ammonium sulfate reduces CBG concentrations” retrieved downregulator with similarity score 0.632, yet was rejected in Stage 3.
These patterns indicate that incorrect rejections stem from overly conservative contextual evaluation during LLM-based refinement in Stage 3 rather than retrieval failure in Stage 2. The system prioritizes precision by applying strict semantic criteria when assessing predicate-context fit, which reduces false positives but increases false rejections in ambiguous cases.
C.3 Negation Handling
We conducted a manual evaluation of negation detection on the HEAL dataset. Of 2,411 relationship extractions from 2,400 abstracts, we identified 123 (5.1%) negated relation texts through manual review.
RELATE correctly flagged 74 as negated assertions, yielding:
  1. 1.
    Precision: 96.1% (74 true positives / 77 flagged as negated)
  2. 2.
    Recall: 60.2% (74 true positives / 123 actual negations)
  3. 3.
    F1-score: 74.2%
The high precision indicates that when RELATE identifies a negation, it is almost always correct (only 3 false positives). The moderate recall indicates that approximately 40% of negated statements are missed (49 false negatives), typically involving:
  1. 1.
    Implicit negations (e.g., “failed to affect”, “unable to deliver”)
  2. 2.
    Comparative negations (e.g., “no significant difference”)
  3. 3.
    Complex syntactic constructions not captured by pattern matching
The negation categories in the 123 manually identified cases were:
  1. 1.
    Explicit negations (“not”, “no”, “does not”): 90 cases (73%)
  2. 2.
    Implicit negations (“failed to”, “unable to”): 21 cases (17%)
  3. 3.
    Comparative negations (“no significant difference”): 12 cases (10%)
RELATE’s negation handling maintains conceptual accuracy by identifying the underlying relationship type while flagging the negation polarity, enabling accurate representation of negative findings in KGs.
Appendix D Predicate and Descriptor Distributions
ChemProt defines 10 classes for chemical-protein interactions: 9 positive relationship types and 1 negative class (NOT/no interaction). We utilize all 9 positive relationship types, yielding 30 positive descriptors (Table LABEL:tab:top_predicates_distribution_chemprot). ChemProt’s NOT class, which indicates no interaction exists, is handled through RELATE’s explicit rejection (NONE) mechanism rather than treating NOT/no interaction as a mappable predicate. This design ensures that only valid biomedical relationships are standardized to ontology predicates while filtering non-relationships before KG integration.
After the preprocessing stage (Section 3.1), the ontology database stores both positive predicates and descriptors and their LLM-negated versions. For ChemProt, this yields 18 stored predicates (9 positive + 9 NEG variants) with 60 total descriptor texts (30 positive + 30 negated). For Biolink, this yields 568 stored predicates (284 positive + 284 NEG variants) with 2,208 total descriptor texts (1,104 positive + 1,104 negated).
Tables LABEL:tab:top_predicates_distribution_chemprot and LABEL:tab:top_predicates_distribution_heal show the descriptor distribution for the most frequent predicates. In ChemProt (Table LABEL:tab:top_predicates_distribution_chemprot), substrate and agonist together account for over one-third of all positive descriptors (36.67%). In Biolink (Table LABEL:tab:top_predicates_distribution_heal), the top 10 predicates account for 41.30% of all positive descriptors, with coexists with (16.12%) and related to (8.70%) being the most prevalent. The remaining 274 Biolink predicates account for 58.70% of descriptors, demonstrating the diversity of biomedical relationship types captured in the ontology.
@@ Appendix D Predicate and Descriptor Distributions
tab:top_predicates_distribution_chemprot Predicate Sample Descriptors Descriptor count (%) substrate Product of, Substrate product of 6 (10.00%) agonist Agonist-activator, Agonist-inhibitor 5 (8.33%) upregulator Activator, Indirect upregulator 4 (6.67%) downregulator Inhibitor, Indirect downregulator 4 (6.67%) regulator Direct regulator, Indirect regulator 3 (5.00%) part of Indicates that the chemical is a structural component or subunit of the protein complex 2 (3.33%) antagonist Chemical substance that binds to and blocks the activation of certain receptors on cells, preventing a biological response 2 (3.33%) modulator Modulator-activator, Modulator-inhibitor 2 (3.33%) cofactor A non-protein chemical compound that is required for an enzyme’s biological activity 2 (3.33%)
Table 3: ChemProt Dataset. All 9 positive relationship predicates ordered by descriptor count. ChemProt’s 10th class (NOT/no interaction) is handled via RELATE’s rejection mechanism (NONE option, Stage 3) rather than predicate mapping.
@@
tab:top_predicates_distribution_heal Predicate Sample Descriptors Descriptor count (%) coexists with holds between two entities that are co-located in the same aggregate object, process, or spatio-temporal region 178 (8.06%) related to holds between two entities 96 (4.39%) located in holds between a material entity and a material entity or site within which it is located (but of which it is not considered a part) 40 (1.81%) part of holds between parts and wholes (material entities or processes) 34 (1.49%) temporally related to holds between two entities with a temporal relationship 30 (1.36%) has part holds between wholes and their parts (material entities or processes) 21 (0.95%) precedes holds between two processes, where one completes before the other begins 15 (0.68%) causes holds between two entities where the occurrence, existence, or activity of one causes the occurrence or generation of the other 15 (0.68%) affects Describes an entity that has an effect on the state or quality of another existing entity 14 (0.63%) has output holds between a process and a continuant, where the continuant is an output of the process 13 (0.63%)
Table 4: Biolink Dataset. Top 10 of 284 predicates by descriptor count. Percentages calculated from 1,104 total positive descriptors across all 284 Biolink predicates used in this study.